# First-passage properties of the jump process with a drift. The general case

### v1

@@ 5.3.1. Divergence of the <<correlation length>>
Let us now study the properties of $S_{T}(\tau\,|\,X_{0})$, the probability that the process remains positive up to time $\tau$. Specifically, we focus the analytic properties of the Laplace transform of the survival probability
@@
in the $\rho$-plane, from which we extract the large $\tau$ behavior of $S_{T}(\tau\,|\,X_{0})$ stated in Sec. 2.1.
@@
The subsequent derivation follows along the lines of Sec. 5. That is, we look at the integrand (125), move the contour of integration from the real line to $\mathcal{C}_{*}(\rho)$ defined by $\mathrm{i}k=\zeta_{*}(\rho)$ and use the resulting expression for the analytic continuation. The main difference from the continuation of $\hat{S}_{N}(s\,|\,\lambda)$ is that now the new contour $\mathcal{C}_{*}(\rho)$ depends on $\rho$, hence, it moves when $\rho$ changes. This gives rise to more technically involved derivations although the general strategy remains the same. For this reason, below we will be concentrating mostly on the details that differ from $\hat{S}_{N}(s\,|\,\lambda)$.
@@
In the following, we separately analyze three cases: the survival regime $\langle M-\alpha t\rangle>0$, the absorption regime $\langle M-\alpha t\rangle<0$, and the critical point $\langle M-\alpha t\rangle=0$.
@@ 6.1. Survival regime
Let us first assume that $\mathrm{Re}\!\left(\lambda+\tfrac{\rho}{\alpha}\right)>0$. In this case according to (55) we have
@@
Now we need to determine whether the deformed contour $\mathcal{C}_{*}(\rho)$ lies above or below the real axis, or, equivalently, to establish the sign of $\zeta_{*}(\rho)$. As argued in Sec. 3.1, this is fixed by the sign of $\mu_{1}(\rho)$. Recall from (29) that
@@
The analytic structure of the integrand in the $k$-plane is similar to that of (82); the logarithm is analytic within the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ and has branch cuts starting at $k=\mathrm{i}\zeta_{1,2}(\rho)$, where $\zeta_{1,2}(\rho)\equiv\zeta_{1,2}(\rho,1)$ are real solutions of (39) with $s=1$ (see Figs. 4 and 5). The term in parentheses provides two simple poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. The schematic representation of the analytic structure is given in Fig. 10.
@@
The case $\mathrm{Re}(\lambda+\frac{\rho}{\alpha})<0$ is treated similarly. The difference is that: (i) the analytic continuation in (126) includes a prefactor; (ii) the contour shift from the real line to $\mathcal{C}_{*}(\rho)$ picks up a residue from the pole at $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. These two contributions cancel, so that the analytic structure in the $\rho$-plane remains unchanged.
Finally, at some point $\rho_{0}$ defined by $\mathrm{Re}(\lambda+\frac{\rho_{0}}{\alpha})>\zeta_{*}(\rho_{0})$, the pole at $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$ traverses the contour $\mathcal{C}_{*}(\rho)$ giving rise to an additional prefactor in front of the integral. As a result, for $\rho_{*}<\rho<\tilde{\rho}$ the analytic continuation of $\hat{S}_{T}(\rho\,|\,\lambda)$ reads
@@
Recall the analogous continuation (88) and (89).
@@
Formally inverting the Laplace transform with respect to $\lambda$ we obtain
@@
Choosing $\gamma$ to be sufficiently large, say $\gamma>-\zeta_{*}(\rho_{*})-\frac{\rho_{*}}{\alpha}$, we see that singularities of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the $\rho$-plane are a pole at $\rho=0$ and the branch point at $\rho=\rho_{*}$.
@@ 6.2. Absorption regime
On the level of analytic structure in the $\rho$-plane, the difference with respect to the survival regime is that $\mu_{1}(0)<0$, and hence $\mathcal{C}_{*}(0)$ lies in the lower part of the $k$-plane. Consequently, when shifting the contour, there is no contribution from the pole at $k=\mathrm{i}\epsilon$, and thus
@@ 6.3. Critical point
We conclude the analysis of the survival probability $S_{T}(\tau\,|\,X_{0})$ by examining the critical point where $\langle M-\alpha t\rangle=0$. The analysis follows closely the approach presented in Sec. 5.3 for $S_{N}(n\,|\,X_{0})$. Therefore, we outline only the main steps and present the key results.
@@ 6.3.1. Divergence of the <<correlation length>>
which according to (137) translates into the divergence
@@
Recalling that $\alpha_{c}=\frac{\langle M\rangle}{\langle t\rangle}$ we obtain the result stated in (18).
@@
From (149) it is clear that $\xi_{\tau}(\alpha_{c})=\infty$ and hence there is no exponential decay of the survival probability $S_{T}(\tau\,|\,X_{0})$, and algebraic behavior is expected instead. On the level of analytic structure of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the $\rho$-plane, it corresponds to the coincidence of the branch point and the pole (recall that at the critical point $\rho_{*}=0$). The large $\tau$ behavior of $S_{T}(\tau\,|\,X_{0})$ is therefore governed by the expansion of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the vicinity of $\rho=0$.
@@
Note that (150) is very similar to (107), and the square root divergence implies that the large $\tau$ behavior is
@@
where $U(X_{0})$ is the same function as for the survival probability $S_{N}(n\,|\,X_{0})$. Recall that it is defined through the Laplace transform as in (109), and the asymptotic behaviors for $X_{0}\to\infty$ and $X_{0}\to 0$ are given by (110) and (111) respectively.
As argued in Sec. 5.3, at the critical point, after a proper rescaling the process converges to a Brownian motion, and the survival probability $S_{N}(n\,|\,X_{0})$ admits a scaling form defined by (118). Repeating essentially the same derivation we obtain
@@ 7. Moments of $\tau$ and $n$
Having established and thoroughly quantified the large $\tau$ and large $n$ behavior of the survival probabilities using the representation (8), the next natural step is to consider other quantities of interest. One important question concerns the moments of the first-passage time $\tau$ and the number of jumps before the first passage $n$ in the absorption regime. In this section we address this question and compute the first two leading terms in the $X_{0}\to\infty$ and $X_{0}\to 0$ expansions for the means and the variances of $\tau$ and $n$.
@@
The representations (156) and (157) however still involve rather complicated integrals, and their practical utility might not be immediately apparent. Especially since we are dealing with a very general class of distributions, and it is unrealistic to expect that the Laplace transforms can be inverted analytically for arbitrary $p(t)$ and $q(M)$.
What can be done instead is to extract the asymptotic behavior of the mean values as $X_{0}\to\infty$ and $X_{0}\to 0$ from the expansions of the Laplace transforms as $\lambda\to 0$ and $\lambda\to\infty$. Typically such expansions can be constructed by expanding the integrand in $I(\lambda;0,1)$ term by term, and the rest becomes a technical exercise. In our case, however, this naive approach results in divergent integrals, necessitating a more careful treatment.
The resolution lies in recognizing that $I(\lambda;\rho,s)$ can be expressed as Mellin convolution integrals, for which powerful asymptotic techniques exist. In this section, we illustrate how such asymptotic expansions can be constructed via Mellin transform techniques, specifically applying the method presented in [63]. We compute the first two terms in the expansion of both the mean and the variance for $\tau$ and $n$ as $X_{0}\to\infty$ and $X_{0}\to 0$ (16 terms in total). This demonstrates that the mapping to an effective random walk, together with the representation (8), provides a powerful tool.
Before proceeding to the technical details, let us make a useful observation. From (156) and (157), the means of $\tau$ and $n$ are related via
@@
Relation (159) can readily be used to recover the asymptotic expansions for the mean value of $\tau$ from those of the mean value of $n$. In other words, instead of 16 terms we need to compute only 12.
@@
Computing the $X_{0}\to\infty$ and $X_{0}\to 0$ asymptotic behaviors of the first two moments involves asymptotic expansions of $I(\lambda;\rho,s)$ (for means) and its first derivatives with respect to $\rho$ and $s$ (for second moments of $\tau$ and $n$) as $\lambda\to 0$ and $\lambda\to\infty$. As previously mentioned, simply expanding the integrand in series with respect to $\lambda$ results in divergent integrals. Leaving the verification of this fact to the reader, we present an alternative approach based on Mellin transform techniques.
@@
which are Mellin convolution integrals. There exists a well-developed body of theory regarding the asymptotic analysis of such integrals. Here we employ the method from [63] for its simplicity and directness.
The key result is as follows. Suppose we aim to find $\lambda\to 0$ behavior of (162), where the functions $f(k)$ and $h(k)$ have asymptotic expansions
@@
or its analytic continuation in the region of $z$ where the integral does not exist.
@@
Furthermore, if $\alpha_{n}-\beta_{m}=1$, then $O\left(\lambda^{\beta_{m}}+\lambda^{\alpha_{n}-1}\right)$ must be changed to $O\left(\lambda^{\beta_{m}}\log\lambda\right)$.
The beauty of (164) together with (167) is that it systematically handles the divergences that would appear if naively expanding (158), converting them into well-defined Mellin transforms.
The final step is to apply (164) to $I(\lambda;0,1)$ and to its derivatives $\partial_{s}I(\lambda;0,1)$ and $\partial_{\rho}I(\lambda;0,1)$. The $\lambda\to 0$ expansions yield the $X_{0}\to\infty$ behavior, while $\lambda\to\infty$ expansions (obtained by reversing the roles of $f$ and $h$) give the $X_{0}\to 0$ behavior. Although the calculations involve considerable algebraic effort, they reduce to systematic application of Taylor expansions and Mellin transform evaluations.
@@
Having established a systematic method to compute the asymptotic expansions of $I(\lambda;0,1)$ and its derivatives, the remaining task is computational. Rather than burden the reader with extensive algebraic details, we defer the computational part to Appendix A and proceed directly to presenting the final results in the form of asymptotic expansions
@@
$a_{0}=\frac{\sqrt{\frac{1}{2}\langle(M-\alpha t)^{2}\rangle}}{\langle M-\alpha t\rangle}e^{-J_{1}},\qquad a_{1}=a_{0}\;L_{1}.$ (182)
@@
$\qquad+a_{0}\frac{\langle M^{2}\rangle\left[\langle t^{2}\rangle-\langle t\rangle^{2}\right]+\langle t^{2}\rangle\left[\langle M^{2}\rangle-\langle M\rangle^{2}\right]}{\langle M-\alpha t\rangle^{2}},$

### v2

@@ 6.3.1. Divergence of the <<correlation length>>
6.4. Numerical verification
To test the validity of the asymptotic results for $S_{N}(n\,|\,X_{0})$, we first verify that they reduce to those of [undefal] when both $p(t)$ and $q(M)$ are exponential distributions; we leave this calculation as an exercise for the reader. To confirm that the results also hold for non-Poissonian arrivals, we perform Monte Carlo simulations.
According to (103) and (106), if $\alpha\neq\alpha_{c}$, then the large-$n$ behavior of the survival probability reads
$\alpha\neq\alpha_{c}:\qquad S_{N}(n\,|\,X_{0})-S_{\infty}(X_{0})\underset{n\to\infty}{\asymp}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right],$ (132)
where the rate $\xi_{n}(\alpha)$ is given by (104), and $S_{\infty}(X_{0})=0$ for $\alpha>\alpha_{c}$. For numerical simulations, to eliminate the constant offset $S_{\infty}(X_{0})$, we analyze the marginal first-passage distribution $\mathbb{P}_{c}[n\,|\,X_{0}]$ conditioned on trajectories in which first passage has occurred:
$\mathbb{P}_{c}[n\,|\,X_{0}]\equiv\frac{1}{1-S_{\infty}(X_{0})}\int_{0}^{\infty}\mathrm{d}\tilde{\tau}\,\mathbb{P}[\tilde{\tau},n\,|\,X_{0}]=\frac{S_{N}(n-1\,|\,X_{0})-S_{N}(n\,|\,X_{0})}{1-S_{\infty}(X_{0})}.$ (133)
In the absorption regime, where the first passage is sure to happen, this is exactly the same as the marginal probability distribution of $n$.
From (132) we expect the exponential decay
$\mathbb{P}_{c}[n\,|\,X_{0}]\underset{n\to\infty}{\asymp}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right].$ (134)
This asymptotic behavior can be refined further. Recall that the function $F(k;0)$ has a saddle point at $k=\mathrm{i}\zeta_{*}(0)$ (Section 4.1), which is precisely where the branch points coincide at $s=s_{*}$. The branch points $\zeta_{1,2}(0,s)$ are determined by the equation $1-sF(\mathrm{i}\zeta;0)=0$. Expanding $F(\mathrm{i}k;0)$ near the saddle point $k=\zeta_{*}$ yields a quadratic form, so that $\zeta_{1,2}(0,s)$ satisfy a quadratic equation in $(\zeta-\zeta_{*})$. They thus exhibit square-root behavior analogous to $A(s)$ and $B(s)$ in the toy example (30) of Section 3. This reveals that the singularity at $s_{*}$ is a square-root branch point, implying
$S_{N}(n\,|\,X_{0})-S_{\infty}(X_{0})\underset{n\to\infty}{\sim}\frac{1}{\sqrt{n}}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right].$ (135)
At the level of the marginal probability, this gives
$\mathbb{P}_{c}[n\,|\,X_{0}]\underset{n\to\infty}{\sim}\frac{1}{n^{3/2}}\exp\left[-\frac{n}{\xi_{n}(\alpha)}\right].$ (136)
To verify the asymptotic behavior of $S_{N}(n\,|\,X_{0})$, we compare analytical predictions with numerical simulations. As in Section 5, we take both $p(t)$ and $q(M)$ to be inverse-Gaussian distributions given by (81) and (82). In this case, the decay rate (104) can be evaluated explicitly. A straightforward computation shows that
$\frac{1}{\xi_{n}(\alpha)}=\left(\ell_{t}\sqrt{\alpha}+\frac{\ell_{M}}{\sqrt{\alpha}}\right)\sqrt{\frac{1}{\ell_{M}+\alpha\ell_{t}}\left(\alpha\frac{\ell_{M}}{m_{M}^{2}}+\frac{\ell_{t}}{m_{t}^{2}}\right)}-\frac{\ell_{M}}{m_{M}}-\frac{\ell_{t}}{m_{t}}.$ (137)
In the numerical simulations, we chose the parameters to be
$m_{t}=1,\qquad\ell_{t}=5,\qquad m_{M}=2,\qquad\ell_{M}=15,\qquad\alpha_{c}=\frac{m_{M}}{m_{t}}=2.$ (138)
With these parameters, we generated $10^{9}$ independent trajectories starting at $X_{0}=0.5$ for two values of the drift velocity: $\alpha=1.6$ and $\alpha=2.4$.
Figure 12 confirms both the exponential decay rate (137) (left column) and the $n^{-3/2}$ power law (right column) through comparison with numerical simulations in both the survival (top row, $\alpha=1.6$) and absorption (bottom row, $\alpha=2.4$) regimes. While plotting the histogram, we retained only values of $n$ that appeared in at least $500$ configurations to ensure adequate statistics.
Figure 12. Comparison of analytical predictions (red dashed lines) with direct Monte Carlo simulations (black circles, $10^{9}$ trajectory realizations for $X_{0}=0.5$) in the survival regime (top row, $\alpha=1.6$) and absorption regime (bottom row, $\alpha=2.4$). Parameters of the distribution are given by (138). Left column: marginal first-passage probability $\mathbb{P}_{c}[n\,|\,X_{0}]$ versus $n$ (logarithmic scale), confirming the exponential decay (134) with rate (137). The theoretical curves are shifted vertically by a factor of $10^{-4}$ for visual comparison. Right column: rescaled probability $e^{n/\xi_{n}(\alpha)}\mathbb{P}_{c}[n\,|\,X_{0}]$ versus $n$ (log-log scale), confirming the $n^{-3/2}$ power law in (136).
At the critical point $\alpha=\alpha_{c}$, the exponential decay (135) is replaced by the power-law behavior, which is given by (117):
$S_{N}(n\,|\,X_{0})\underset{n\to\infty}{\sim}\frac{1}{\sqrt{\pi n}}U(X_{0}).$ (139)
Unlike the rates in the off-critical regimes, the asymptotic expansions (119) and (120) of $U(X_{0})$ for $X_{0}\to\infty$ and $X_{0}\to 0$ require numerical integration. For the parameters (138) we obtain
$U(X_{0})\underset{X_{0}\to 0}{\approx}1.06-1.03\;X_{0},\qquad U(X_{0})\underset{X_{0}\to\infty}{\approx}1.22\;X_{0}+0.90.$ (140)
We computed $S_{N}(n\,|\,X_{0})$ for $n=100$ numerically across a range of initial positions $X_{0}$ to test three distinct regimes.
Figure 13. Comparison of analytical predictions (dashed and dotted lines) with direct Monte Carlo simulations (black circles, $10^{6}$ trajectory realizations for each value of $X_{0}$) for $S_{N}(n=100\,|\,X_{0})$ at the critical point $\alpha=\alpha_{c}$. The parameters of the distributions are given by (138). The left panel shows moderate initial positions $X_{0}\in[0,3]$, where the data are described by large-$n$ behavior as in (139). The blue dotted and red dashed lines represents the small-$X_{0}$ and large-$X_{0}$ expansions (140) respectively. The right panel extends to larger $X_{0}$ values, displaying the crossover to the universal scaling regime. The blue dotted line shows the scaling function (127), and the red dashed line again shows the large-$X_{0}$ expansion. The gray vertical line at $X_{0}=\sqrt{n}$ indicates the crossover scale.
The results shown in Fig. 13 demonstrate agreement between analytical predictions and Monte Carlo simulations. The left panel focuses on moderate initial positions $X_{0}\in[0,3]$ where $n=100$ is sufficiently large for the asymptotic behavior (139) to apply. For $X_{0}\lesssim 1$, the small-$X_{0}$ expansion (blue dotted line) accurately captures the numerical data. For $X_{0}\gtrsim 1$, the large-$X_{0}$ expansion (red dashed line) provides a description instead. The right panel extends to larger initial positions. In this regime, the asymptotic limit $n\to\infty$ with fixed $X_{0}$ no longer applies. Instead, the data approach the universal scaling limit (127) (blue dotted line), which depends only on the ratio $X_{0}/\sqrt{n}$. The crossover between the large-$X_{0}$ asymptotic and the scaling regime occurs near $X_{0}\approx\sqrt{n}$ (gray vertical line), as predicted.
@@ 6.4. Numerical verification
Let us now study the properties of $S_{T}(\tau\,|\,X_{0})$, the probability that the process remains positive up to time $\tau$. Specifically, we focus on the analytic properties of the Laplace transform of the survival probability
@@
in the $\rho$-plane, from which we extract the large-$\tau$ behavior of $S_{T}(\tau\,|\,X_{0})$ stated in Section 2.1.
@@
The subsequent derivation follows along the lines of Section 6. That is, we examine the integrand (143), move the contour of integration from the real line to $\mathcal{C}_{*}(\rho)$ defined by $\mathrm{i}k=\zeta_{*}(\rho)$ and use the resulting expression for the analytic continuation. The main difference from the continuation of $\hat{S}_{N}(s\,|\,\lambda)$ is that now the new contour $\mathcal{C}_{*}(\rho)$ depends on $\rho$, hence, it moves when $\rho$ changes. This gives rise to more technically involved derivations, although the general strategy remains the same. For this reason, below we be concentrate mostly on the details that differ from $\hat{S}_{N}(s\,|\,\lambda)$.
@@
In the following, we separately analyze three cases: the survival regime $\langle M-\alpha t\rangle>0$, the absorption regime $\langle M-\alpha t\rangle<0$, and the critical point $\langle M-\alpha t\rangle=0$. Finally, we compare the analytical results with the numerical simulations.
@@ 7.1. Survival regime
Let us first assume that $\mathrm{Re}\!\left(\lambda+\tfrac{\rho}{\alpha}\right)>0$. In this case, according to (59), we have
@@
Now we need to determine whether the deformed contour $\mathcal{C}_{*}(\rho)$ lies above or below the real axis, or, equivalently, to establish the sign of $\zeta_{*}(\rho)$. As argued in Section 4.1, this is fixed by the sign of $\mu_{1}(\rho)$. Recall from (45) that
@@
The analytic structure of the integrand in the $k$-plane is similar to that of (91); the logarithm is analytic within the strip $\mathrm{Im}(k)\in(\zeta_{1},\zeta_{2})$ and has branch cuts emanating from $k=\mathrm{i}\zeta_{1,2}(\rho)$, where $\zeta_{1,2}(\rho)\equiv\zeta_{1,2}(\rho,1)$ are real solutions of (55) with $s=1$ (see Figs. 6 and 7). The term in parentheses provides two simple poles at $k=\mathrm{i}\epsilon$ and $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. The schematic representation of the analytic structure is given in Fig. 14.
@@
The case $\mathrm{Re}(\lambda+\frac{\rho}{\alpha})<0$ is treated similarly. The differences are: (i) the analytic continuation in (144) includes a prefactor; (ii) the contour shift from the real line to $\mathcal{C}_{*}(\rho)$ picks up a residue from the pole at $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$. These two contributions cancel, so that the analytic structure in the $\rho$-plane remains unchanged.
Finally, at some point $\rho_{0}$ defined by $\mathrm{Re}(\lambda+\frac{\rho_{0}}{\alpha})=-\zeta_{*}(\rho_{0})$, the pole at $k=-\mathrm{i}(\lambda+\frac{\rho}{\alpha})$ traverses the contour $\mathcal{C}_{*}(\rho)$ giving rise to an additional prefactor in front of the integral. As a result, for $\rho_{*}<\rho<\tilde{\rho}$ the analytic continuation of $\hat{S}_{T}(\rho\,|\,\lambda)$ reads
@@
Recall the analogous expressions (97) and (98) for $\hat{S}_{N}(s\,|\,\lambda)$.
@@
Formally inverting the Laplace transform with respect to $\lambda$, we obtain
@@
Choosing $\gamma$ to be sufficiently large, say $\gamma>-\zeta_{*}(\rho_{*})-\frac{\rho_{*}}{\alpha}$, we see that the singularities of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the $\rho$-plane are a pole at $\rho=0$ and the branch point at $\rho=\rho_{*}$.
@@ 7.2. Absorption regime
At the level of analytic structure in the $\rho$-plane, the difference from the survival regime is that $\mu_{1}(0)<0$, and hence $\mathcal{C}_{*}(0)$ lies in the lower part of the $k$-plane. Consequently, when shifting the contour, there is no contribution from the pole at $k=\mathrm{i}\epsilon$, and thus
@@ 7.3. Critical point
We conclude the analysis of the survival probability $S_{T}(\tau\,|\,X_{0})$ by examining the critical point where $\langle M-\alpha t\rangle=0$. The analysis closely follows the approach presented in Section 6.3 for $S_{N}(n\,|\,X_{0})$. Therefore, we outline only the main steps and present the key results.
@@ 7.3.1. Divergence of the <<correlation length>>
which, according to (155), translates into the divergence
@@
Recalling that $\alpha_{c}=\frac{\langle M\rangle}{\langle t\rangle}$, we obtain the result stated in (19).
@@
From (167) it is clear that $\xi_{\tau}(\alpha_{c})=\infty$ and hence there is no exponential decay of the survival probability $S_{T}(\tau\,|\,X_{0})$; algebraic behavior is expected instead. At the level of the analytic structure of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the $\rho$-plane, it corresponds to the coincidence of the branch point and the pole (recall that at the critical point $\rho_{*}=0$). The large $\tau$ behavior of $S_{T}(\tau\,|\,X_{0})$ is therefore governed by the expansion of $\hat{S}_{T}(\rho\,|\,\lambda)$ in the vicinity of $\rho=0$.
@@
Note that (168) is very similar to (116), and the square root divergence implies that the large-$\tau$ behavior is
@@
where $U(X_{0})$ is the same function as for the survival probability $S_{N}(n\,|\,X_{0})$. Recall that it is defined through the Laplace transform as in (118), and the asymptotic behaviors for $X_{0}\to\infty$ and $X_{0}\to 0$ are given by (119) and (120), respectively.
As argued in Section 6.3, at the critical point, after a proper rescaling the process converges to a Brownian motion, and the survival probability $S_{N}(n\,|\,X_{0})$ admits a scaling form defined by (127). Repeating essentially the same derivation, we obtain
@@
Figure 15. Comparison of analytical predictions (red dashed lines) with direct Monte Carlo simulations (black circles, $10^{9}$ trajectory realizations for $X_{0}=0.5$) in the survival regime (top row, $\alpha=1.6$) and absorption regime (bottom row, $\alpha=2.4$). Parameters of the distribution are given by (138). Left column: marginal first-passage probability $\mathbb{P}_{c}[\tau\,|\,X_{0}]$ versus $\tau$ (logarithmic scale), confirming the exponential decay (173) with rate (174). The theoretical curves are shifted vertically by a factor of $10^{-4}$ for visual comparison. Right column: rescaled probability $e^{\tau/\xi_{\tau}(\alpha)}\mathbb{P}_{c}[\tau\,|\,X_{0}]$ versus $\tau$ (log-log scale), confirming the $\tau^{-3/2}$ power law in (173).
7.4. Numerical verification
To verify the asymptotic behavior of the survival probability $S_{T}(\tau\,|\,X_{0})$, we use the same simulations as in Section 6.4 for $S_{N}(n\,|\,X_{0})$. Both distributions $p(t)$ and $q(M)$ are inverse-Gaussian as given by (81) and (82) with parameters (138). The only difference is that now we concentrate on the marginal first-passage time probability distribution $\mathbb{P}_{c}[\tau\,|\,X_{0}]$ conditioned on the trajectories in which the first passage occurs:
$\mathbb{P}_{c}[\tau\,|\,X_{0}]=-\frac{1}{1-S_{\infty}(X_{0})}\frac{\partial}{\partial\tau}\Big[S_{T}(\tau\,|\,X_{0})\Big].$ (172)
Repeating the arguments of Section 6.4 we expect the right tail of the distribution to be given by
$\alpha\neq\alpha_{c}:\qquad\mathbb{P}_{c}\left[\tau\,|\,X_{0}\right]\underset{\tau\to\infty}{\sim}\frac{1}{\tau^{3/2}}\exp\left[-\frac{\tau}{\xi_{\tau}(\alpha)}\right]$ (173)
Substituting (84) and (83) into (155), a straightforward computation shows that
$\frac{1}{\xi_{\tau}(\alpha)}=\frac{1}{2}\frac{\ell_{M}\ell_{t}}{\ell_{M}+\alpha\ell_{t}}\left(\frac{1}{m_{t}}-\frac{\alpha}{m_{M}}\right)^{2}.$ (174)
Figure 15 demonstrates that the behavior (173) is correct. The results of the simulations at the critical point $\alpha=\alpha_{c}$ are shown in Fig. 16.
Figure 16. Comparison of analytical predictions (dashed and dotted lines) with direct Monte Carlo simulations (black circles, $10^{6}$ trajectory realizations for each value of $X_{0}$) for $S_{T}(\tau=150\,|\,X_{0})$ at the critical point $\alpha=\alpha_{c}$. The parameters of the distributions are given by (138). The left panel shows moderate initial positions $X_{0}\in[0,3]$, where the data are described by large-$\tau$ behavior as in (170). The blue dotted and red dashed lines represents the small-$X_{0}$ and large-$X_{0}$ expansions (140), respectively. The right panel extends to larger $X_{0}$ values, displaying the crossover to the universal scaling regime. The blue dotted line shows the scaling function (171), and the red dashed line again shows the large-$X_{0}$ expansion. The gray vertical line at $X_{0}=\sqrt{\tau}$ indicates the crossover scale.
@@ 8. Moments of $\tau$ and $n$
Having established and thoroughly quantified the large-$\tau$ and large-$n$ behavior of the survival probabilities using the representation (8), the next natural step is to consider other quantities of interest. In the absorption regime, an important question concerns the moments of the first-passage time $\tau$ and the number of jumps $n$ before the first passage. In this section, we address this question and compute the first two leading terms in the $X_{0}\to\infty$ and $X_{0}\to 0$ expansions for the means and the variances of $\tau$ and $n$.
@@
The representations (177) and (178), however, still involve rather complicated integrals, and their practical utility might not be immediately apparent. This is especially true since we are dealing with a very general class of distributions, and it is unrealistic to expect that the Laplace transforms can be inverted analytically for arbitrary $p(t)$ and $q(M)$.
What we can do instead is to extract the asymptotic behavior of the mean values as $X_{0}\to\infty$ and $X_{0}\to 0$ from the expansions of the Laplace transforms as $\lambda\to 0$ and $\lambda\to\infty$. Typically such expansions can be constructed by expanding the integrand in $I(\lambda;0,1)$ term by term, and the rest becomes a technical exercise. In our case, however, this naive approach results in divergent integrals, necessitating a more careful treatment.
The resolution lies in recognizing that $I(\lambda;\rho,s)$ can be expressed as a Mellin convolution integral, for which powerful asymptotic techniques exist. In this section, we illustrate how such asymptotic expansions can be constructed via Mellin transform techniques, using the method presented in [undefaan]. We compute the first two terms in the expansion of both the mean and the variance for $\tau$ and $n$ as $X_{0}\to\infty$ and $X_{0}\to 0$ (16 terms in total). This demonstrates that the mapping to an effective random walk, together with the representation (8), provides a powerful tool.
Before proceeding to the technical details, let us make a useful observation. From (177) and (178), it is clear that the means of $\tau$ and $n$ are related via
@@
This relation can readily be used to recover the asymptotic expansions for the mean value of $\tau$ from those of the mean value of $n$. In other words, instead of 16 terms, we need to compute only 12.
@@
Computing the $X_{0}\to\infty$ and $X_{0}\to 0$ asymptotic behaviors of the first two moments involves asymptotic expansions of $I(\lambda;\rho,s)$ (for means) and its first derivatives with respect to $\rho$ and $s$ (for second moments of $\tau$ and $n$) as $\lambda\to 0$ and $\lambda\to\infty$. As previously mentioned, simply expanding the integrand in series with respect to $\lambda$ results in divergent integrals. We leave verification of this fact to the reader and instead present an alternative approach based on Mellin transform techniques.
@@
which are Mellin convolution integrals. There exists a well-developed body of theory regarding the asymptotic analysis of such integrals. Here we employ the method from [undefaan] for its simplicity and directness.
The key result is as follows. Suppose we aim to find the $\lambda\to 0$ behavior of (183), where the functions $f(k)$ and $h(k)$ have asymptotic expansions
@@
or, when integral diverges, its analytic continuation.
@@
Furthermore, if $\alpha_{n}-\beta_{m}=1$, then the error term $O\left(\lambda^{\beta_{m}}+\lambda^{\alpha_{n}-1}\right)$ becomes $O\left(\lambda^{\beta_{m}}\log\lambda\right)$.
The beauty of (185) together with (188) is that it systematically handles the divergences that would appear from naive expansion of (179), converting them into well-defined Mellin transforms.
The final step is to apply (185) to $I(\lambda;0,1)$ and to its derivatives $\partial_{s}I(\lambda;0,1)$ and $\partial_{\rho}I(\lambda;0,1)$. The $\lambda\to 0$ expansions yield the $X_{0}\to\infty$ behavior, while the $\lambda\to\infty$ expansions (obtained by reversing the roles of $f$ and $h$) give the $X_{0}\to 0$ behavior. Although the calculations involve considerable effort, they reduce to systematic application of Taylor expansions and Mellin transform evaluations.
@@
Now that we have established a method to compute the asymptotic expansions of $I(\lambda;0,1)$ and its derivatives, the remaining task is computational. Rather than burden the reader with extensive algebraic details, we defer the computational part to Appendix A and proceed directly to presenting the final results in the form of asymptotic expansions
@@
$a_{0}=-\frac{\sqrt{\frac{1}{2}\langle(M-\alpha t)^{2}\rangle}}{\langle M-\alpha t\rangle}e^{-J_{1}},\qquad a_{1}=-a_{0}\;L_{1}.$ (204)
@@
$\qquad+a_{0}\frac{\langle M\rangle^{2}\left[\langle t^{2}\rangle-\langle t\rangle^{2}\right]+\langle t\rangle^{2}\left[\langle M^{2}\rangle-\langle M\rangle^{2}\right]}{\langle M-\alpha t\rangle^{2}},$